The market demand curve for microwaves is Q = 260 - 3P.


The market supply curve for microwaves is Q = 20 + P.


The equilibrium price of a microwave is 60 (Enter only a number. Do NOT enter any symbols.)


The equilibrium quantity of microwaves is 80 (Enter only a number. Do NOT enter any symbols.)


Suppose the government gives firms a $20 subsidy for every microwave sold. After the subsidy is imposed:


The price consumers pay for a microwave after each microwave is subsidized is  (Enter only a number. Do NOT enter any symbols.)


The price firms receive for a microwave after each microwave is subsidized is   (Enter only a number. Do NOT enter any symbols.)


The quantity of microwaves exchanged in the market after each microwave is subsidized is   (Enter only a number. Do NOT enter any symbols.)




1
Expert's answer
2021-07-05T16:18:59-0400

The equilibrium point is when the demand and supply curves are equal.

Hence, "260 - 3P = 20 + P"

"240 = 4P"

"P = 60"

Substituting we get "Q = 80"

"1. 60 2 \times 80"

After subsidy, the price actually paid by the consumer is"(P-20)" . Assuming the supply is constant the lower price paid by the consumer will increase demand thus increasing the equilibrium price and quantity.

"260 - 3(P - 20) = 20 + P"

"260 - 3P + 60 = 20 + P"

"P = 75"

Substituting we get "Q = 95 3 \times (75-20) = 52415"